# Which chromium-sulfur compounds exist as 2D material?

### v1

@@ Abstract
Two-dimensional (2D) chromium-sulfides are synthesized by molecular beam epitaxy using graphene as a substrate. Structure characterization by employing scanning tunneling microscopy and low energy electron diffraction indicates that there are two 2D phases, Cr2S3-2D and CrS4-2D, which have not been reported before. CrS4-2D is related to bulk Cr5S6, but thinner than a bulk unit cell. For Cr2S3-2D, an even thinner material, no bulk counterpart exists. Both 2D materials are found to be structurally stable under ambient conditions and exhibit interesting electronic properties. Extensive first-principles calculations provide further insight into the electronic structure of these systems and indicate that they should be magnetic. Surprisingly, CrS2 recently reported from density functional theory calculations and experiments could not be fabricated.
1 Introduction
@@
Research on chromium-chalcogene compound 2D materials has recently flourished, with significant advancements both theoretically in the prediction of phases and properties as well as experimentally in the preparation and investigation of new phases 1, 2, 3. One element of excitement, shared with chromium halides 4, 5, 6, is the interesting magnetic properties of these new 2D materials 7, 8, 9, 10, 11, 12, 13, 14, 15. Focusing on the case of chromium-sulfur 2D materials 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41 it seems not yet to be settled which phases exist as 2D materials. Neither were all compounds predicted from DFT calculations experimentally fabricated nor is the experimental research comprehensive enough to make additional discoveries unlikely. This situation is one of the motivations for the present study.
@@
While the structure and composition of 2D materials created by exfoliation are known from their bulk crystal, the situation is far more intricate for 2D materials grown by molecular beam epitaxy (MBE) or chemical vapor deposition (CVD). These 2D materials may possess no bulk counterpart; thus, their structure and composition are not known a priori. The thinness and lack of signal make standard techniques like X-ray diffraction or Raman spectroscopy difficult to apply. The unknown arrangement of planes of different species, depth-dependent attenuation, and vacancy-containing planes impair even composition analysis methods like X-ray photoelectron spectroscopy (XPS). Frequently, density functional theory (DFT) calculations are used as a guideline for interpreting experimental compounds, although stability in DFT calculations does neither guarantee a pathway for creation nor the existence of a specific phase. All of this is relevant to 2D materials based on chromium and sulfur.
@@
The bulk chromium sulfur phase diagram is complex 42, 43, 44, 45, but no Cr-rich phases have been documented. Unquestioned S-rich phases are limited in stoichiometry by Cr2S3, except of the more S-rich Cr5S8 phase which can be prepared under high pressure. Noteworthy, no compound with van der Waals gaps has been reported as bulk phase.
@@
Bulk Cr2S3 can be considered to be composed of hypothetical CrS2 layers with octahedral (T) Cr coordination, where the interlayer gaps are filled by 1/3 of the possible sites establishing chemical bonds between the CrS2 layers. Depending on the precise conditions of preparation and composition, the distribution of interlayer Cr atoms may be disordered or give rise to trigonal or rhombohedral stacking [see Figure 1(a)] 43, 46, 47. For the latter two, the arrangement of the interlayer Cr repeats every two or three CrS2 layers, respectively. Complex magnetically ordered structures are reported, which critically depend on the arrangement of Cr interlayer atoms and the exact composition 46, 48, 49, 47, 50. The magnetic moments are in-plane, normal to the c-axis 48, 47. Rhombohedral Cr2S3 is a semiconductor, paramagnetic above and ferrimagnetic below 120 K with a maximum magnetization around 90 K 49, 46, 47, 50. Also negative magnetostriction and photoconductivity were reported for rhombohedral Cr2S3 50, 51.
@@
Up to now, only the rhombohedral variant of Cr2S3 has been synthesized as 2D material 16, 17, 18, 19, 20, 21, 22, 23, 24, using either chemical vapor deposition or sulfurization of a metallic film. Single unit cell thickness has been reported 16, 17, 19, 22, 24, with the unit cell consisting of four CrS2 layers with three planes of 1/3 Cr sites occupied in the interlayer gaps. However, in a transmission electron microscopy study the arrangement of Cr interlayer atoms in as-grown "rhombohedral" CVD films was found to be random 22. Similar to bulk counterpart, semiconducting behavior, magnetization with a T K, negative magnetostriction, and good photosensitivity were found 17, 20, 19.
@@
While no bulk compound of composition CrS2 has been reported, DFT calculations proposed the existence of CrS2 as a typical transition metal dichalcogenide with the S atomic planes sandwiching the Cr atomic plane [see Figure 1(b)] 37, 36. H- (trigonal prismatic Cr coordination), T-, and T’- phases were identified as possible variants, with ambiguity on the most stable phase being either the H- 37, 31 or the T-phase 36. Recently, also two different orthorhombic CrS2 phases were identified in calculations 33, 52. DFT based research for CrS2 materials is flourishing due to exciting predicted properties ranging from ferromagnetism 36, 33, 31 with Curie temperatures up to 1000 K, over a direct band gap with valley polarization 37 to excellent performance for the hydrogen evolution reaction 35, to name a few.
@@
Despite the absence of the bulk compound, the synthesis of CrS2 as 2D material has been communicated25, 30, 27, 26, 28, in the T-phase 25, 27, 26, 28, 40, 39, but also in the H- and T’-phases 25. CrS2 islands tend to grow thicker than single unit cell, but in two cases minimal heights in the range of nm were measured 25, 28. Transport in CrS2 field effect devices was characterized as p-type 25, n-type 26, or metallic 27, 28. Weak ferromagnetic behavior up to room temperature 27 and use as a floating gate in photoelectric non-volatile memory 28 were described.
@@
In bulk Cr5S6, 2/3 of the possible Cr sites in the interlayer gaps are filled, i.e., it is a material in the NiAs structure with 1/3 of Cr vacancies in every second Cr layer [see Figure 1(c)] 43. Cr5S6 is antiferromagnetic below 168 K and ferrimagnetic up to 303 K 48. Up to now, to the authors’ best knowledge, preparation of Cr5S6 as a 2D material has not yet been reported.
@@
Lastly, there is a yet unconfirmed DFT prediction for a new CrxSy 2D material, namely Cr2S3-2D [see Figure 1(d)] 40, 39. To avoid confusion with bulk materials with the same composition but different structure, "-2D" is attached to the stoichiometric formulas in the 2D case from hereon. Cr2S3-2D is unrelated to bulk Cr2S3 and consists of five planes of atoms in the sequence S-Cr-S-Cr-S in the NiAs structure, which can be related to bulk CrS, although it lacks the monoclinic distortion observed there 43. Based on the DFT calculations, Cr2S3-2D is proposed to be a magnetic semiconductor with a transition temperature above room temperature (405 K) 53. The key difference of Cr2S3-2D compared to other 2D magnetic semiconductors, such as CrI3, is the presence of both intralayer and interlayer exchange coupling 40. This characteristic of Cr2S3-2D may account for its high transition temperature.
@@
2 Results and discussion
2.1 Structure from experiment
CrxSy-2D compounds are grown on single crystal graphene (Gr) prepared on Ir(110). For the LEED pattern in Figure 2(a) after deposition of Cr in S vapor at room temperature, the sample was post-annealed in S vapor at 750 K (see Methods for details). Besides the Gr/Ir(110) substrate reflections 54, a diffraction ring is also visible. It has a sixfold symmetric intensity modulation with the highest intensity aligned with the first-order Gr reflections. We assign this diffraction ring with radius to the newly formed CrxSy-2D compound. Based on the sixfold symmetry of the diffraction ring, a threefold or sixfold in-plane symmetry of the new compound can be assumed, resulting in an in-plane lattice constant of nm.
The typical topography of the sample is presented through the STM image of Figure 2(b). Gr/Ir(110) is covered with an area fraction of about by coalesced CrxSy-2D islands with wavy edges and typical linear dimensions of 50 nm. Faint dark lines are visible within the islands, which we interpret as grain boundaries formed during annealing due to the coalescence of almost randomly oriented smaller islands. They are best visible in the contrast-enhanced box in Figure 2(b). There, also the larger nm Gr/Ir(110) moiré periodicity can be recognized which is imprinted from the substrate onto the island. The atomically-resolved STM image of CrxSy-2D in the inset displays a surface lattice of periodicity of nm consistent with determined from LEED. As shown by the height profile along the white line in Figure 2(b), the islands display a uniform apparent height of nm at a sample bias V. Irrespective of the precise preparation conditions, we never observed islands of lower height. As justified below, we interpret this CrxSy-2D compound as Cr2S3-2D of which the side view ball model is shown in Figure 1(d).
@@ 2.1 Structure from experiment
While phase pure Cr2S3-2D is formed up to 800 K, upon increasing the annealing temperature, the transformation of Cr2S3-2D is not limited to CrS4-2D, but thicker and thicker layers are formed [see Figure S1 in the Supporting Information (SI)]. In the absence of an S vapor during annealing or for larger initial deposited amounts, the transformation of Cr2S3-2D to thicker layers sets in already at lower temperatures (see Figure S2 and Figure S3 of the SI). While the CrxSy-2D materials thicker than CrS4-2D are beyond the scope of the present manuscript, we point out that the apparent heights for V increase in the sequence 0.79 nm, 1.02 nm, 1.29 nm, 1.56 nm, 1.82 nm and so forth (see Figure S3), i.e., the levels above Cr2S3-2D increase in steps of 0.23 nm to 0.27 nm.
@@
Considering that Cr2S3-2D contains 2 Cr atoms per unit cell as established above, by taking into account the deposition times and coverages measured, we obtain a Cr content of for CrS4-2D. The experimentally found Cr content is thus consistent with CrS4-2D containing or 2.67 Cr atoms per unit cell.
@@
Based on the experimental information and considering that CrxSy-2D phases are alternating sequences of S and partial or complete Cr planes 43, we are ready to make an informed hypothesis about the structure of the two CrxSy-2D phases. For Cr2S3-2D we assume a structure as depicted in Figure 1(d). It contains 2 Cr atoms per unit cell, as found experimentally. In principle, one S plane sandwiched between two Cr planes would be consistent with our experiments. However, the reactivity of the metal and growth under sulfur-rich conditions excludes the presence of bare metal at the Gr interface (see ref. 55) or the vacuum. Therefore, we assume the presence of two additional sulfur planes, making the compound inert and resulting in the stoichiometry Cr2S3-2D. Although we cannot guess the precise stacking, it is plausible to be similar to the bulk compounds Cr2S3 or Cr5S6 [Figure 1(a) and (c)], when disregarding the presence of Cr vacancies in these compounds. The observed hexagonal surface symmetry with a lattice parameter of 0.341 nm matches reasonably with the lattice parameter of the Cr deficient bulk compounds Cr2S3 and Cr5S6, which is in the range from 0.343-0.345 nm. Previous DFT calculations for Cr2S3-2D yield lattice parameters of 0.343-0.348 nm 39, 40 in acceptable agreement with our experiment. We emphasize that the absence of any superstructure together with 2 Cr atoms per unit cell is strong evidence for complete Cr planes in Cr2S3-2D.
@@
An additional partial or complete Cr plane plus a saturating S plane has for bulk CrxSy materials an average height of 0.28 nm 43, in close agreement with the step height of 0.26-0.27 nm measured for thicker CrxSy-2D by us. The slightly smaller apparent step height between Cr2S3-2D and CrS4-2D of 0.23 nm fits reasonably well, and the slight deviation may be of electronic origin. It is, therefore, most plausible to assume now the presence of an additional 2/3 Cr plane (based on the measured additional Cr amount of 0.62 of a full plane) and an additional S plane giving rise to the proposed stoichiometry CrS4-2D. The presence of a 2/3 Cr plane within the stack provides a natural explanation for the observed -R30∘ superstructure, although the position of the 2/3 Cr plane is uncertain. Based on the surface sensitivity of LEED signals, it is unlikely that the 2/3 Cr plane is the lowermost in the stack. Whether the partial Cr plane is in the center as indicated in Figure 1(e) or at the top cannot be decided based on the experimental data provided up to now.
@@
Lastly, the stability and inertness of the covalently bound Cr2S3-2D is highlighted through Figure 5. As documented by the two subsequent STM topographs of Figure 5(a) and (b), a large Cr2S3-2D island can be laterally shifted by the STM-tip interacting with the island edge without significant defect formation. This is a direct proof of just a weak van der Waals interaction between Cr2S3-2D and Gr/Ir(111). Figure 5(c) displays an STM topograph of Cr2S3-2D islands taken in a vacuum after exposure for three hours to ambient conditions [see Figure S6 for a large-scale topography of the same sample]. Not unexpectedly, adsorption on Gr and Cr2S3-2D is apparent, but the structural integrity of the islands is obvious.
2.2 Absence of CrS2

### v2

@@ Abstract
Two-dimensional (2D) chromium-sulfides are synthesized by molecular beam epitaxy using graphene as a substrate. Structure characterization by employing scanning tunneling microscopy and low energy electron diffraction indicates that there are two 2D phases, Cr2S3-2D and CrS4-2D, which have not been reported before. CrS4-2D is related to bulk Cr5S6, but thinner than a bulk unit cell. For Cr2S3-2D, an even thinner material, no bulk counterpart exists. Both 2D materials are found to be structurally stable under ambient conditions and exhibit interesting electronic properties. Extensive first-principles calculations provide further insight into the electronic structure of these systems and indicate that they should be magnetic. Although single layers of CrS2 were predicted to be stable by density functional theory calculations and reported in previous experimental studies, we were unable to synthesize CrS2 under our range of experimental conditions.
I Introduction
@@
Research on chromium-chalcogene compound 2D materials has recently flourished, with significant advancements both theoretically in the prediction of phases and properties as well as experimentally in the preparation and investigation of new phases [1, 2, 3]. One element of excitement, shared with chromium halides [4, 5, 6], is the interesting magnetic properties of these new 2D materials [7, 8, 9, 10, 11, 12, 13, 14, 15]. Focusing on the case of chromium-sulfur 2D materials [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41] it seems not yet to be settled which phases exist as 2D materials. Neither were all compounds predicted from DFT calculations experimentally fabricated nor is the experimental research comprehensive enough to make additional discoveries unlikely. This situation is one of the motivations for the present study.
@@
While the structure and composition of 2D materials created by exfoliation are known from their bulk crystal, the situation is far more intricate for 2D materials grown by molecular beam epitaxy (MBE) or chemical vapor deposition (CVD). These 2D materials may possess no bulk counterpart; thus, their structure and composition are not known a priori. The thinness and lack of signal make standard techniques like X-ray diffraction or Raman spectroscopy difficult to apply. The unknown arrangement of planes of different species, depth-dependent attenuation, and vacancy-containing planes impair even composition analysis methods like X-ray photoelectron spectroscopy (XPS). Frequently, density functional theory (DFT) calculations are used as a guideline for interpreting experimental compounds, although stability in DFT calculations guarantees neither a pathway for creation nor the existence of a specific phase. All of this is relevant to 2D materials based on chromium and sulfur.
@@
The bulk chromium sulfur phase diagram is complex [42, 43, 44, 45], but no Cr-rich phases have been documented. Unquestioned S-rich phases are limited in stoichiometry by Cr2S3, except of the more S-rich Cr5S8 phase which can be prepared under high pressure. Noteworthy, no bulk phase with van der Waals gaps has been reported.
@@
Bulk Cr2S3 can be considered to be composed of hypothetical CrS2 layers with octahedral (T) Cr coordination, where the interlayer gaps are filled by 1/3 of the possible sites establishing chemical bonds between the CrS2 layers. Depending on the precise conditions of preparation and composition, the distribution of interlayer Cr atoms may be disordered or give rise to trigonal or rhombohedral stacking [see Figure 1(a)] [43, 46, 47]. For the latter two, the arrangement of the interlayer Cr repeats every two or three CrS2 layers, respectively. Complex magnetically ordered structures are reported, which critically depend on the arrangement of Cr interlayer atoms and the exact composition [46, 48, 49, 47, 50]. The magnetic moments are in-plane, normal to the c-axis [48, 47]. Rhombohedral Cr2S3 is a semiconductor, paramagnetic above and ferrimagnetic below 120 K with a maximum magnetization around 90 K [49, 46, 47, 50]. Negative magnetostriction and photoconductivity were also reported for rhombohedral Cr2S3 [50, 51].
@@
Rhombohedral [16, 17, 18, 19, 20, 21, 22, 23, 24] and trigonal [52] Cr2S3 were synthesized as thin films,either using chemical vapor deposition or sulfurization of a metallic film. 2D material of few unit cell thickness for trigonal [52] and single unit cell thickness for rhombohedral Cr2S3 [16, 17, 19, 22, 24] was reported. For the latter, the unit cell consists of four CrS2 layers with three planes of 1/3 Cr sites occupied in the interlayer gaps, just as shown in Figure 1(a)]. However, in a transmission electron microscopy study the arrangement of Cr interlayer atoms in as-grown ”rhombohedral” CVD films was found to be random [22]. Similar to the bulk counterpart, semiconducting behavior, magnetization with a T K, negative magnetostriction, and good photosensitivity were found [17, 20, 19].
@@
No bulk crystals of composition CrS2 have yet been reported. The closest to the existence of bulk CrS2 are misfit layer compounds. These crystals consist of layers of monoclinic CrS2 alternating with layers of LaS, BiS or Y [53, 54, 55]. Although bulk CrS2 is absent, DFT calculations proposed the existence of CrS2 as a typical transition metal dichalcogenide with the S atomic planes sandwiching the Cr atomic plane [see Figure 1(b)] [37, 36]. H- (trigonal prismatic Cr coordination), T-, and T’- phases were identified as possible variants, with ambiguity on the most stable phase being either the H- [37, 31] or the T-phase [36]. Recently, also two different orthorhombic CrS2 phases were identified in calculations [33, 56]. DFT based research for CrS2 materials is flourishing due to exciting predicted properties ranging from ferromagnetism [36, 33, 31] with Curie temperatures up to 1000 K, over a direct band gap with valley polarization [37] to excellent performance for the hydrogen evolution reaction [35], to name a few.
@@
Despite the absence of the bulk compound, the synthesis of CrS2 as 2D material has been communicated [25, 30, 27, 26, 28], in the T-phase [25, 27, 26, 28, 40, 39], but also in the H- and T’-phases [25]. CrS2 islands tend to grow thicker than single unit cell, but in two cases minimal heights in the range of nm were measured [25, 28]. Transport in CrS2 field effect devices was characterized as p-type [25], n-type [26], or metallic [27, 28]. Weak ferromagnetic behavior up to room temperature [27] and use as a floating gate in photoelectric non-volatile memory [28] were described.
@@
In bulk Cr5S6, 2/3 of the possible Cr sites in the interlayer gaps are filled, i.e., it is a material in the NiAs structure with 1/3 of Cr vacancies in every second Cr layer [see Figure 1(c)] [43]. Cr5S6 is antiferromagnetic below 168 K and ferrimagnetic up to 303 K [48]. Until now, to the authors’ best knowledge, preparation of Cr5S6 as a 2D material has not yet been reported.
@@
Lastly, there is a yet unconfirmed DFT prediction for a new CrxSy 2D material, namely Cr2S3-2D [see Figure 1(d)] [40, 39]. To avoid confusion with bulk materials with the same composition but different structure, ”-2D” is attached to the stoichiometric formulas in the 2D case from hereon. Cr2S3-2D is unrelated to bulk Cr2S3 and consists of five planes of atoms in the sequence S-Cr-S-Cr-S in the NiAs structure, which can be related to bulk CrS, although it lacks the monoclinic distortion observed there [43]. Based on DFT calculations, Cr2S3-2D is proposed to be a magnetic semiconductor with a transition temperature above room temperature (405 K) [57]. The key difference of Cr2S3-2D compared to other 2D magnetic semiconductors, such as CrI3 (ordering temperature K [4]) and CrSBr (ordering temperature K [58]), is the presence of both intralayer and interlayer exchange coupling [40]. The additional presence of interlayer exchange coupling might be a decisive factor enhancing the magnetic ordering temperature of Cr2S3-2D. Increasing covalency of bonds is an additional factor enhancing superexchange interactions [59]. Therefore one may expect an increasing ordering temperature from CrI3 over CrSBr to Cr2S3-2D.
@@
II Results and discussion
II.1 Structure from experiment
CrxSy-2D compounds are grown on graphene (Gr) prepared on low index Ir single crystal surfaces. Gr is a suitable inert van der Waals substrate that enables the facile growth of transition-metal sulfide compounds, leaving their intrinsic properties largely unaffected [60]. On Ir(111) and Ir(110) Gr can be grown prior to CrxSy-2D under ultrahigh vacuum (UHV) conditions in single crystal quality [61, 62]. Thus, all measurements shown here result from samples prepared under UHV conditions that have never left the UHV environment, and are consequently ultimately clean. The presentation here will focus on the results using the Gr/Ir(110) substrate. The results for Gr/Ir(111) show only minute differences, as will be mentioned in more detail on the sideline below.
For the LEED pattern in Figure 2(a) after deposition of Cr in S vapor at room temperature, the sample was post-annealed in S vapor at 750 K (see Methods for details). In addition to the Gr/Ir(110) substrate reflections [62], a diffraction ring is also visible. It has a sixfold symmetric intensity modulation with the highest intensity aligned with the first-order Gr reflections. We assign this diffraction ring with radius to the newly formed CrxSy-2D compound. Based on the sixfold symmetry of the diffraction ring, a threefold or sixfold in-plane symmetry of the new compound can be assumed, resulting in an in-plane lattice constant of nm.
The typical topography of the sample is presented through the STM image of Figure 2(b). Gr/Ir(110) is covered with an area fraction of about by coalesced CrxSy-2D islands with wavy edges and typical linear dimensions of 50 nm. Faint dark lines are visible within the islands, which we interpret as grain boundaries formed during annealing because of the coalescence of almost randomly oriented smaller islands. They are best visible in the contrast-enhanced box in Figure 2(b). There, also the larger nm Gr/Ir(110) moiré periodicity can be recognized, which is imprinted from the substrate onto the island. The atomically resolved STM image of CrxSy-2D in the inset displays a surface lattice of periodicity of nm consistent with determined from LEED. As shown by the height profile along the white line in Figure 2(b), the islands display a uniform apparent height of nm at a sample bias V. Regardless of the precise preparation conditions, we never observed islands of lower height. As justified below, we interpret this CrxSy-2D compound as Cr2S3-2D of which the side view ball model is shown in Figure 1(d).
@@ II.1 Structure from experiment
While phase pure Cr2S3-2D is formed up to 800 K, upon increasing the annealing temperature, the transformation of Cr2S3-2D is not limited to CrS4-2D, but thicker and thicker layers are formed [see Figure S2, Supporting Information]. In the absence of an S vapor during annealing or for larger initial deposited amounts, the transformation of Cr2S3-2D to thicker layers sets in already at lower temperatures (see Figure S3, Supporting Information). While the CrxSy-2D materials thicker than CrS4-2D are beyond the scope of the present manuscript, we point out that the apparent heights for V increase in the sequence 0.79 nm, 1.02 nm, 1.29 nm, 1.56 nm, 1.82 nm and so forth (see Figure S4, Supporting Information), i.e., the levels above Cr2S3-2D increase in steps of 0.23 nm to 0.27 nm.
@@
Considering that Cr2S3-2D contains 2 Cr atoms per unit cell as established above, by taking into account the deposition times and coverages measured, we obtain a Cr content of for CrS4-2D. The experimentally found Cr content is thus consistent with CrS4-2D containing or 2.67 Cr atoms per original unit cell or 8 Cr atoms for the superstructure unit cell.
@@
Based on the experimental information and considering that CrxSy-2D phases are alternating sequences of S and partial or complete Cr planes [43], we are ready to make an informed hypothesis about the structure of the two CrxSy-2D phases. For Cr2S3-2D we assume a structure as depicted in Figure 1(d). It contains 2 Cr atoms per unit cell, as found experimentally. In principle, one S plane sandwiched between two Cr planes would be consistent with our experiments. However, the reactivity of the metal and growth under sulfur-rich conditions excludes the presence of bare metal at the Gr interface (see ref. 63) or the vacuum. Therefore, we assume the presence of two additional sulfur planes, making the compound inert and resulting in the stoichiometry Cr2S3-2D. Although we cannot guess the precise stacking, it is plausible to be similar to the bulk compounds Cr2S3 or Cr5S6 [Figure 1(a) and (c)], when disregarding the presence of Cr vacancies in these compounds. The observed hexagonal surface symmetry with a lattice parameter of 0.341 nm matches reasonably with the lattice parameter of the Cr deficient bulk compounds Cr2S3 and Cr5S6, which is in the range from 0.343-0.345 nm. Previous DFT calculations for Cr2S3-2D yield lattice parameters of 0.343-0.348 nm [39, 40] in acceptable agreement with our experiment. We emphasize that the absence of any superstructure together with 2 Cr atoms per unit cell is strong evidence for complete Cr planes in Cr2S3-2D.
@@
An additional partial or complete Cr plane plus a saturating S plane has for bulk CrxSy materials an average height of 0.28 nm [43], in close agreement with the step height of 0.26-0.27 nm measured for thicker CrxSy-2D by us. The slightly smaller apparent step height between Cr2S3-2D and CrS4-2D of 0.23 nm fits reasonably well, and the slight deviation may be of electronic origin. It is, therefore, most plausible to assume now the presence of an additional 2/3 Cr plane (based on the measured additional Cr amount of 0.62 of a full plane) and an additional S plane giving rise to the proposed stoichiometry CrS4-2D. The presence of a 2/3 Cr plane within the stack provides a natural explanation for the observed -R30∘ superstructure, although the position of the 2/3 Cr plane is uncertain. Based on the surface sensitivity of LEED signals, it is unlikely that the 2/3 Cr plane is the lowermost in the stack. Whether the partial Cr plane is in the center as indicated in Figure 1(e) or at the top cannot be decided based on the experimental data provided up to now.
@@
Lastly, the stability and inertness of the covalently bound Cr2S3-2D is highlighted through Figure 5. As documented by the two subsequent STM topographs of Figure 5(a) and (b), a large Cr2S3-2D island can be laterally shifted by the STM-tip interacting with the island edge without significant defect formation. This is a direct proof of a weak van der Waals interaction between Cr2S3-2D and Gr/Ir(111). Figure 5(c) displays an STM topograph of Cr2S3-2D islands taken in vacuum after exposure for three hours to ambient conditions [see Figure S5, Supporting Information, for a before and after exposure comparison and a large-scale topography of the same sample]. Not unexpectedly, adsorption on Gr and Cr2S3-2D is apparent, but the structural integrity of the islands is obvious.
II.2 Absence of CrS2